OpenStudy (anonymous):

A 0.16 kg billiard ball moving at 1.20 m/s collides elastically with one of the same mass at rest. It continues at a speed of 0.60 m/s afterward in a different direction from before. What is the final speed of the other billiard ball?

OpenStudy (anonymous):

I hope I can finish this question quickly! I have to go to work soon.

OpenStudy (ybarrap):

This is a completely elastic collision, so by the conservation of momentum: $$ m_1v_1=mv_1=m_1v_1'+m_2v_2'=m(v_1'+v_2')\\ v_2'=v_1-v_1'\\ =1.2-0.60 $$

OpenStudy (anonymous):

So it's just 0.60 m/s?

OpenStudy (ybarrap):

yes

OpenStudy (anonymous):

You are amazing :D haha! Thank you so much for helping me today!!

OpenStudy (ybarrap):

you're welcome!